# Beyond Coordinates: Meta-Equivariance in Statistical Inference

### v1

@@ 4.4 Implications
Acknowledgements
This paper emerged from an accelerated process of independent discovery—developed over four days, and refined briefly thereafter. But its pace is not the story. Its process is.
What you’re reading is a proof-of-concept in what theoretical research becomes when a single human mind is extended through live coordination with high-resolution computational tools. The result is not just a paper, but an instance of a new mode of mathematical cognition—one where structure, simulation, and symbolic form co-evolve in real time.
I’m a second-year undergraduate in economics. I developed this work independently, without formal supervision, and with no prior background in the relevant mathematical fields. That isn’t a disclaimer—it’s a statement of the underlying method.
The central results—including the principle of meta-equivariance—originated not from traditional derivation, but from a symbiotic workflow where ideas emerged via interaction between human intuition and machine verification. Hypotheses were tested numerically before they were formalised. Patterns appeared in Python before they were proved on paper. Formal structure followed empirical signal.
This approach is sometimes called centaur cognition: a tightly coupled collaboration between human insight and algorithmic precision. It’s not outsourcing to AI, nor is it human-in-the-loop micromanagement. It’s a real-time feedback system in which intuition guides computation, computation sharpens intuition, and the boundary between idea and execution collapses. The AI in this process functioned less as a tool and more as an interface for latent structure—mirroring, constraining, challenging, and accelerating the search for form.
This is one of eight papers developed in similarly compressed cognitive cycles over the last two months. Each is backed by rigorous code pipelines, numerical validation to machine epsilon, and self-contained formal derivations. But more importantly, each is a test case for this emerging method—a methodology where the bottleneck is not proof-writing, but the bandwidth of insight.
This paper stands in two registers: first, as a contribution to statistical theory; second, as a prototype of an alternative scholarly workflow. A system where insight and verification move at the same speed. Where formal training matters—but is no longer the gatekeeper of discovery.
I welcome all feedback, scrutiny, and correction. Meta-equivariance is only one result among several. My goal is not just to share conclusions—but to share the method itself. If this process can work for me—an undergraduate with no institutional guidance—it could transform what’s possible for those already trained in the language of mathematics.
If you’re reading this and you want to test the method, collaborate, or dismantle it—I would like to hear from you. My email is open to anyone interested in or skeptical of this workflow.

### v2

@@ 4.4 Implications
Author’s Note
I am a second-year undergraduate in economics at the University of Bristol working independently. This paper belongs to a research programme I am developing across information theory, econometrics and mathematical statistics.
This work was produced primarily through AI systems that I directed and orchestrated. The AI generated the mathematical content, proofs and symbolic derivations based on my research questions and guidance. I have no formal mathematical training but am eager to learn through this process of directing AI-powered mathematical exploration.
My contribution involves designing research directions, evaluating and selecting AI outputs, and ensuring the coherence of the overall research agenda. All previously published work that influenced these results has been cited to the best of my knowledge and research capabilities in my current position. The presentation aims to be pedagogically accessible.